OpenStudy (anonymous):

Use basic identities to simplify the expression. (cos of theta^2 / sin of theta ^2) + csc θ sin θ

Nnesha (nnesha):

\[\huge\rm \frac{ \cos^2 \theta }{ \sin ^2 \theta } +\csc \theta \sin \theta \] like this ?

Nnesha (nnesha):

write csc in terms of sin or cos what's the definition of csc ?? like sec = 1/cos csc = ??

OpenStudy (anonymous):

csc = 1/sin theta would I multiply sin theta * 1/sin theta and get 1?

Nnesha (nnesha):

yes that's right :-)

Nnesha (nnesha):

and cos /sin = what ? :-) remember tan and cot are reciprocal of each other

OpenStudy (anonymous):

(1/ sec theta)^2 /(1 / csc theta)^2?? I'm not really sure how to simplify that :(

Nnesha (nnesha):

well that's right but simple way is to change cos/sin to cot \[\tan =\frac{ \sin }{ \cos } ~~\cot=\frac{ \cos }{ \sin }\]

Nnesha (nnesha):

so cos^2 theta over sin ^2 theta = cot^2 theta

OpenStudy (anonymous):

is there a way to simplify that further?

Nnesha (nnesha):

yes

Nnesha (nnesha):

\[\huge\rm\color{blue}{ \frac{ \cos^2 \theta }{ \sin ^2 \theta } }+\color{reD}{\csc \theta \sin \theta} \] \[\large\rm \color{blue}{ \cot^2 \theta} +\color{red}{1} = ???\] which is equal to what ??it's an identity

OpenStudy (anonymous):

pythagoras? csc^2 theta?

Nnesha (nnesha):

yes right cot^2 + 1 = csc^2 x

OpenStudy (anonymous):

omg thank you! identities confuse me so much :)

Nnesha (nnesha):

my pleasure :-) and practice!! :-)